How do i get rid of the fraction for 2/x - 3/y = 1
Answer by nerdybill(7384)

By multiplying both sides by a "common denominator" -- xy:
.
2/x - 3/y = 1
xy(2/x - 3/y) = xy(1)
2y - 3x = xy